As an epidemiologist you are going to investigate the effect of a drug suspected of causing malformations in newborn infants when the drug in question is taken by pregnant women during the course of their pregnancies. As your sample you will use the next 200 single births occurring in a given hospital. For each birth a medication history will be taken from the new mother and from her doctor; in addition, you will review medical records to verify use of the drug. [N.B.: These mothers are considered to have been followed prospectively during the entire course of their pregnancies, because a complete and accurate record of drug use was maintained during pregnancy.]
The resultant data are:
Forty mothers have taken the suspected drug during their pregnancies. Of these mothers, 35 have delivered malformed infants. In addition, 10 other infants are born with malfunctions.
Suppose that a relative risk of 0.5 has been found. Which interpretation might follow?

A.

The risk of congenital malformations is half as high among users of the drug as among non-users.

B.

The risk of congenital malformations is equal among users and non-users of the drug.

C.

The risk of congenital malformations is twice as high among users of the drug as among non-users of the drug.

D.

Use of the drug is a protective factor for congenital malformations.
Answer #1

relative risk = risk among non users / risk among users

relative risk = 0.5 = 1/2

risk among non users / risk among users = 1/2

risk among non users / risk among users = 1/2

risk among users = 2 * (risk among non users)

answer : option C : The risk of congenital malformations is twice as high among users of the drug as among non-users of the drug.

(please UPVOTE)